# Strong and noise-tolerant entanglement in dissipative optomechanics

### v2

@@ Abstract
Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in optomechanical systems with purely dispersive coupling over the past decades. However, quantum entanglement, induced by purely dissipative coupling, remains unexplored. In this work, we study quantum entanglement in a Michelson–Sagnac interferometer, where the dipersive and the dissipative coupling can be arbitrarily switched on and off. With current experimental parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with the purely dispersive (dissipative) coupling. Further, we find that quantum entanglement generated by purely dissipative coupling is significantly stronger and more noise-tolerant than that generated by purely dispersive coupling. When both couplings coexist, entanglement is weakened due to interference. Our findings provide a promising path to engineer strong and noise-tolerant quantum entanglement in purely dissipative quantum systems.
@@ I Introduction
Due to its potential applications in gravitational wave detection Abbott et al. (2016) and quantum-classical boundary tests Haroche (1998), cavity optomechanics has garnered significant theoretical and experimental interest in quantum information science Aspelmeyer et al. (2014). With advances in ground-state cooling Liu et al. (2013); Teufel et al. (2011); Gröblacher et al. (2009); Schliesser et al. (2008), exploring quantum effects in cavity optomechanics has become a reality Lü et al. (2015); Safavi-Naeini et al. (2013); Purdy et al. (2013); Weis et al. (2010); Safavi-Naeini et al. (2011); Kronwald and Marquardt (2013); Koppenhöfer et al. (2023); Guo et al. (2016); Kyriienko et al. (2014a); Xiong et al. (2016, 2021a); Chen et al. (2021); Lü et al. (2013); Peng et al. (2023); Xiong et al. (2023); Wang et al. (2024); Mann et al. (2018); Jäger et al. (2019); Xiong et al. (2021b, 2022), especially for macroscopic quantum entanglement Vitali et al. (2007). Over the past two decades, numerous theoretical proposals have been put forward to achieve optomechanical entanglement Tian (2012, 2013); Wang and Clerk (2013); Chen et al. (2024, 2023); Liu et al. (2024a); Lai et al. (2022); Jiao et al. (2020, 2022); Liu et al. (2024b); Shang and Li (2024). Experimentally, macroscopic entanglement between two massive oscillators has also been realized in cavity optomechanics experimentally Ockeloen-Korppi et al. (2018). Notably, both theoretical and experimental proposals have predominantly relied on coherent dispersive coupling mechanisms between the cavity mode and the mechanical mode.
@@
Very recently, unconventional dissipative optomechanics was first theoretically proposed in a superconducting microwave circuit Elste et al. (2009) and later extended to a Michelson-Sagnac interferometers (MSI) with a movable membrane Xuereb et al. (2011). In such dissipative optomechanics, the cavity linewidth depends on the mechanical displacement Elste et al. (2009), giving rise to the dissipative coupling between the mechanical and cavity modes. This unique coupling mechanism opens a potential avenue to investigate intriguing quantum phenomena, such as unconventional bistability Kyriienko et al. (2014b), quantum squeezing Qu and Agarwal (2015), enhanced ground-state cooling Huang and Chen (2018); Liu et al. (2023), quantum sensing Huang and Agarwal (2017), negative-damping instability Elste et al. (2009), self-sustained oscillations Huang et al. (2018), and phonon lasing Kuang et al. (2023); Zhang et al. (2022). Experimentally, dissipative optomechanics have been demonstrated in various platforms, including the MSI Sawadsky et al. (2015); Friedrich et al. (2011), microdisc resonator coupled to a nanomechanical waveguide Li et al. (2009), photonic crystal split-beam nanocavity Wu et al. (2014), graphene drums coupled to high-Q microsphere Cole et al. (2015), and levitated sphere trapped by dual-beam optical tweezers Kuang et al. (2023). It is evident that dissipative optomechanics has made significant progress both theoretically and experimentally. However, how dissipative optomechanics influences quantum entanglement, particularly under dissipative coupling or purely dissipative coupling conditions, remains an open and unexplored question.
@@
In this work, we propose to investigate quantum entanglement in an experimentally feasible dissipative optomechanical system Sawadsky et al. (2015); Friedrich et al. (2011). The setup consists of a fixed perfect mirror and an effective movable mirror, realized by a Michelson–Sagnac interferometer (MSI) incorporating a membrane. Compared to dispersive optomechanics Vitali et al. (2007) with negligible dissipative coupling, the dispersive and dissipative couplings can be independently switched on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter (BS) in this unconventional dissipative optomechanics Xuereb et al. (2011). We first examine the behavior of the steady-state mechanical displacement with purely dispersive or dissipative coupling. It is evident that that the displacement exhibits nonlinear dependence on the driving power with purely dispersive coupling, similar to conventional optomechanical systems Aspelmeyer et al. (2014). However, it becomes linear dependence of the driving power for purely dissipative coupling. We then study the effect of the purely dispersive or dissipative coupling on generation of quantum entanglement. We surprisely find that the quantum entanglement induced by purely dissipative coupling is not only much stronger than that generated by purely dispersive coupling, but also it is more robust aginst the bath temperature. For purely dispersive entanglement, its survival temperature is approximately $\sim 8.5\,\text{K}$, while for purely dissipative entanglement, it can be up to $\sim 25\,\text{K}$. When both couplings coexist, entanglement reduces due to quantum interference. Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with purely dissipative coupling.
@@ III Steady-state solution
When dissipation is included, the dynamics of the proposed system can be governed by the quantum Langevin equantion $\mathcal{O}=-i[\mathcal{O},\mathcal{H}_{S}]+{Dissipation}+{Noise}$, specifically,
@@
Figure 2: The normalized steady-state displacment $x_{s}$ of the membrane vs (a) the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ and (b) the power of the driving field. The red curve corresponds to the case of the purely dispersive coupling ($g_{\omega}/2\pi=2$ Hz). The blue curve is plotted for the purely dissipative coupling ($g_{\kappa}/2\pi=2$ Hz). In (a) and (b), $\Delta_{a}=3\kappa_{a}$ is assumed.
@@
Obivously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative coupling between photons and phonons. For a purely dispersive coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig. 2(a)]. For a purely dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
@@
where $x_{s}\neq 2\kappa_{a}/g_{\kappa}$ and $n_{s}\leq(\kappa_{a}/g_{\kappa})^{2}(\omega_{m}/2\Delta_{a})$ are both required. When $n_{s}=0$, $x_{s}=0$, i.e., the membrane is positioned at its equilibrum. Here $x_{s}$ and $n_{s}$ are quadratic dependence for the purely dissipative coupling, demonstrated by the blue curve in Fig. 2(a). One may further notice that the displacement of the membrane caused by the purely dispersive coupling is larger than that caused by the purely dissipative coupling, as clearly shown in Fig. 2(a). To plot Fig. 2(a), we adopt typical experimental parameters Sawadsky et al. (2015): the wavelength of the driving field $\lambda=1064$ nm, $\omega_{d}/2\pi=c/\lambda=281.96$ THz, the quality factor of the membrane $\mathcal{Q}_{m}=5.8\times 10^{5}$, $\omega_{m}/2\pi=136$ kHz, $\gamma_{m}/2\pi=0.23$ Hz, $L=8.7$ cm, $m=80$ ng, and $\kappa_{a}/2\pi=1.5$ MHz. These experimental parameters ensure the stability of the proposed system according to Routh-Hurwitz criterion.
@@
Explicitly, Eq. (11) may have three real roots, corresponding to bistability. But in fact, the bistability can only be predicted for puly dispersive coupling, as shown by the red curve in Fig. 2(b). This indicates that the purely dissipative coupling can not generate the typical stability in optomechanics, demonstrated by the blue curve in Fig. 2(b).
@@ IV Covariance matrix
With steady-state values in hand, we further expand each operator $\mathcal{O}$ around its steady state, i.e., $\mathcal{O}=\delta\mathcal{O}+\mathcal{O}_{s}$. Substituting this expression back into Eq. (III), we then obtain a set of equations for fluctuations ($\delta\mathcal{O}$). For the strong driving field ($|a_{s}|^{2}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al. (2007), corresponding to linearization. As a result, the dynamics for fluctuations can be written as
@@
Compared to the case of purely dispersive coupling, the dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, rendering $D$ non-diagonal.
@@ V Optomechanical Entanglement
To investigate optomechanical entanglement, we consider three specific scenarios: (i) purely dispersive coupling, (ii) purely dissipative coupling, and (iii) the coexistence of both couplings.
V.1 Purely dispersive coupling ($g_{\kappa}=0$)
@@
In the purely dispersive coupling regime ($g_{\kappa}=0$), we present the entanglement measure $E_{N}$ as a function of both the dispersive coupling strength $g_{\omega}$ and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$ in Fig. 3(a) to explore the conditions for optimal optomechanical entanglement at a bath temperature of $T=0.4$ K. The results reveal that the generated entanglement remains relatively weak compared to that typically observed in conventional optomechanical systems. This is primarily due to the inherent limitation in the dispersive coupling strength $g_{\omega}$ (or equivalently $G_{\omega}$), which must remain small within the stable regime to avoid system instability, thereby restricting the achievable entanglement. Nonetheless, optimal entanglement can still be achieved around the point ($\Delta_{s}/\omega_{m},g_{\omega}/2\pi)\approx(2,3.1)$, as marked by the green star, indicating that even under these constraints, entanglement can be optimized.
@@
Figure 3: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the purely dispersive coupling strength $g_{\omega}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the dispersive coupling strength $g_{\omega}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\omega}/2\pi=3.1$ Hz and $\Delta_{s}/\omega_{m}=2$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW.
@@
To further investigate the role of dispersive coupling strength on entanglement, we plot $E_{N}$ as a function of $g_{\omega}$ in Fig. 3(b), where $\Delta_{s}\approx 2\omega_{m}$ is fixed. It is obviously shown that $E_{N}$ increases with $g_{\omega}$. Once up to its maximum, $E_{N}$ will have a sharp reduction by further increasing $g_{\omega}$. Such abrupt behavior corresponds to the transition from the stability to instability, as shown by the pink-shaded region. Additionally, when the bath temperature increases to $T=4$ K [see the blue dashed curve], the coupling for generation of quantum entanglement is required to be larger and $E_{N}$ has a significant reduction. This indicates that quantum entanglement can be affected by the bath temperature. To clearly evaluate the effect of bath temperature, we plot $E_{N}$ as a function of the bath temperature in Fig. 3(c), where $g_{\omega}/2\pi=3.1$ Hz is fixed to obtain optimal entanglement. As the temperature increases, $E_{N}$ gradually declines. When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, that is, quantum entanglment vanishes and the system becomes separable.
V.2 Purely dissipative coupling ($g_{\omega}=0$)
Figure 4: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the purely dissipative coupling strength $g_{\kappa}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the dissipative coupling strength $g_{\kappa}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=0.1$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\kappa}/2\pi=19$ Hz and $\Delta_{s}/\omega_{m}=0.1$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW.
@@ V.2 Purely dissipative coupling ($g_{\omega}=0$)
In the purely dissipative case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the purely dispersive case, even surpassing the entanglement achieved in conventional optomechanical systems (i.e., $E_{N}=0.35$). This is demonstrated in Fig. 4(a), where $E_{N}$ is plotted as a function of both the dissipative coupling strength ($g_{\kappa}$) and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$. The optimal entanglement is found at approximately ($\Delta_{s}/\omega_{m},g_{\kappa}/2\pi)\approx(0.1,20)$, as indicated by the green star. In contrast to the purely dispersive case [see Fig. 2(a)], where the optimal entanglement occurs near the mechanical red sideband at $\Delta_{s}/\omega_{m}=2$, the optimal entanglement in the purely dissipative case is obtained when the cavity is nearly resonant with the driving field, specifically at $\Delta_{s}/\omega_{m}=0.1$, which is far removed from the mechanical red sideband.
@@
Furthermore, in Fig. 4(b), we investigate the behavior of the optomechanical entanglement as a function of the dissipative coupling strength, while fixing $\Delta_{s}=0.1\omega_{m}$. The resulting behavior of $E_{N}$ mirrors that observed in the purely dispersive coupling case (see the red curve). The abrupt drop in entanglement marks the transition of the system from a stable to an unstable region, as indicated by the pink area. Notably, a relatively large dissipative coupling strength is required to achieve the optimal entanglement. This suggests that in the regime dominated by dissipative coupling, stronger interactions are necessary to generate significant quantum effects. As the bath temperature increases, for instance to $T=4$ K, the entanglement is reduced (see the blue curve). To explore the influence of bath temperature on optomechanical entanglement, we fix $g_{\kappa}/2\pi=19$ Hz and present the results in Fig. 4(c). We observe a sharp decrease in entanglement as the system temperature increases from 0 K to 5 K. However, beyond this point, the entanglement becomes more tolerant to further temperature increases, suggesting that the dissipative coupling becomes more robust to thermal noise at higher temperatures. Remarkably, entanglement persists even up to 25 K, This indicates that strong, noise-resistant entanglement can be generated through purely dissipative coupling, opening new possibilities for quantum information processing and communication in noisy

### v3

@@ Abstract
Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in coherent optomechanics over the past decades. However, entanglement in dissipative optomechanics, where the cavity linewidth is dependent on the position of the mechanical resonator, remains unexplored. In this work, we study quantum entanglement in a dissipative optomechanics, realized by a Michelson–Sagnac interferometer and a movable membrane. Such dissipative optomechanics allows to switch coherent and dissipative optomechanical coupling at will. With reliable parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with coherent (dissipative) coupling. Further, we show that quantum entanglement generated by dissipative coupling is significantly stronger and more noise-tolerant than that generated by coherent coupling. When both the coherent and dissipative couplings are simultaneously considered, entanglement is instead weakened due to quantum interference. Our result indicates that dissipative optomechanical coupling may be a promising path to engineer strong and noise-tolerant quantum entanglement.
@@ I Introduction
Coherent optomechanics, where the cavity frequency is dependent on the position of the mechanical resonator, has garnered siginificant interest theoretically and experimentally over the past two decades Aspelmeyer et al. (2014). In such systems, various proposals have been put forward, including ground-state cooling Teufel et al. (2011); Gröblacher et al. (2009); Schliesser et al. (2008), bistability Dorsel et al. (1983); Gozzini et al. (1985) and tristability Xiong et al. (2016), optomechancally induced transparency Weis et al. (2010); Safavi-Naeini et al. (2011); Kronwald and Marquardt (2013), quantum squeezing Liao and Law (2011); Nunnenkamp et al. (2010); Lü et al. (2015); Aggarwal et al. (2020); Purdy et al. (2013); Luo et al. (2022), phase transition Wang et al. (2024); Jäger et al. (2019); Mann et al. (2018); Xu et al. (2017), microwave-optical conversion Guo et al. (2016); Han et al. (2021); Sahu et al. (2022); Andrews et al. (2014); Chen et al. (2023a); Xu et al. (2016), spin-optomechanical interface Xiong et al. (2021a); Chen et al. (2021); Peng et al. (2023); Xiong et al. (2023); Jäger et al. (2019), higher-order exceptional point Xiong et al. (2021b, 2022), and macroscopic quantum entanglement Vitali et al. (2007); Tian (2012, 2013); Wang and Clerk (2013); Chen et al. (2024, 2023b); Liu et al. (2024a); Lai et al. (2022); Jiao et al. (2020, 2022); Liu et al. (2024b); Shang and Li (2024); Ockeloen-Korppi et al. (2018). Recently, dissipative optomechanics was theoretically proposed in a superconducting microwave circuit Elste et al. (2009), where the cavity linewidth is a function of the mechanical displacement Elste et al. (2009), giving rise to a dissipative coupling between the mechanical and cavity modes. Later, the setup was extended to a Michelson-Sagnac interferometers (MSI) with a movable membrane Xuereb et al. (2011), where coherent or dissipative optomechanical coupling is allowed to switch on and off at will by tuning the position of the membrane or the reflectivity of the beam splitter. Experimentally, dissipative optomechanics have been demonstrated in MSI Sawadsky et al. (2015); Friedrich et al. (2011), microdisc resonator coupled to a nanomechanical waveguide Li et al. (2009), photonic crystal split-beam nanocavity Wu et al. (2014), graphene drums coupled to high-Q microsphere Cole et al. (2015), and levitated sphere trapped by dual-beam optical tweezers Kuang et al. (2023). Within dissipative optomechanics, diverse intriguing quantum phenomena, such as unconventional bistability Kyriienko et al. (2014), quantum squeezing Qu and Agarwal (2015); Huang and Chen (2020), enhanced ground-state cooling Huang and Chen (2018); Liu et al. (2023), quantum sensing Huang and Agarwal (2017), negative-damping instability Elste et al. (2009), self-sustained oscillations Huang et al. (2018), and phonon lasing Kuang et al. (2023); Zhang et al. (2022) have been investigated. However, macroscopic quantum entanglement has not been revealed yet with dissipative optomechanical coupling, although it has been intensively studied in coherent optomechanics.
@@
In this work, we propose to investigate quantum entanglement in an experimentally feasible dissipative optomechanical system Sawadsky et al. (2015); Friedrich et al. (2011). The setup consists of a fixed perfect mirror and an effectively movable mirror, realized by a Michelson–Sagnac interferometer (MSI) with an integrated membrane. We begin by analyzing the steady-state mechanical displacement under both coherent and dissipative coupling. Notably, with coherent coupling, the mechanical displacement exhibits a nonlinear dependence on the driving power, consistent with conventional optomechanics Aspelmeyer et al. (2014). In contrast, under dissipative coupling, the displacement shows a linear dependence on the driving power. Furthermore, we explore quantum entanglement in the presence of coherent and dissipative coupling. Interestingly, we find that entanglement induced by dissipative coupling is not only significantly stronger than that generated by coherent coupling but also remarkably more robust against bath temperature. For entanglement with coherent optomechanical coupling, its survival temperature is approximately $\sim 8.5\,\text{K}$, while for entanglement with dissipative optomechanical coupling, it can be up to $\sim 25\,\text{K}$. When both the coherent and dissipative couplings are simultaneously considered, entanglement instead reduces due to quantum interference. Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with dissipative coupling.
@@ III Steady-state solution
When dissipation is included, the dynamics of the proposed system can be governed by the quantum Langevin equantion $\dot{\mathcal{O}}=-i[\mathcal{O},\mathcal{H}_{S}]+{Dissipation}+{Noise}$, specifically,
@@
Figure 2: The normalized steady-state displacment $x_{s}$ of the membrane vs (a) the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ and (b) the power of the driving field. The red curve corresponds to the case of the coherent coupling ($g_{\omega}/2\pi=2$ Hz). The blue curve is plotted for the dissipative coupling ($g_{\kappa}/2\pi=2$ Hz). In (a) and (b), $\Delta_{a}=3\kappa_{a}$ is assumed.
@@
Obivously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative coupling between photons and phonons. For a coherent coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig. 2(a)]. For a dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
@@
where $x_{s}\neq 2\kappa_{a}/g_{\kappa}$ and $n_{s}\leq(\kappa_{a}/g_{\kappa})^{2}(\omega_{m}/2\Delta_{a})$ are both required. When $n_{s}=0$, $x_{s}=0$, i.e., the membrane is positioned at its equilibrum. Here $x_{s}$ and $n_{s}$ are quadratic dependence for the dissipative coupling, demonstrated by the blue curve in Fig. 2(a). One may further notice that the displacement of the membrane caused by the coherent coupling is larger than that caused by the dissipative coupling, as clearly shown in Fig. 2(a). To plot Fig. 2(a), we adopt typical experimental parameters Sawadsky et al. (2015): the wavelength of the driving field $\lambda=1064$ nm, $\omega_{d}/2\pi=c/\lambda=281.96$ THz, the quality factor of the membrane $\mathcal{Q}_{m}=5.8\times 10^{5}$, $\omega_{m}/2\pi=136$ kHz, $\gamma_{m}/2\pi=0.23$ Hz, $L=8.7$ cm, $m=80$ ng, and $\kappa_{a}/2\pi=1.5$ MHz. These experimental parameters ensure the stability of the proposed system according to Routh-Hurwitz criterion.
@@
Explicitly, Eq. (11) may have three real roots, corresponding to bistability. But in fact, the bistability can only be predicted for coherent coupling, as shown by the red curve in Fig. 2(b). This indicates that dissipative coupling can not generate the typical bistability in optomechanics, demonstrated by the blue curve in Fig. 2(b).
@@ IV Covariance matrix
With steady-state values in hand, we further expand each operator $\mathcal{O}$ around its steady state, i.e., $\mathcal{O}=\delta\mathcal{O}+\mathcal{O}_{s}$. Substituting this expression back into Eq. (III), we then obtain a set of equations for fluctuations ($\delta\mathcal{O}$). For the strong driving field ($n_{s}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al. (2007), corresponding to linearization. As a result, the dynamics for fluctuations can be written as
@@
Compared to coherent coupling case, dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, playing an crucial role in generation of bipartite entanglement below.
@@ V Optomechanical Entanglement
To investigate optomechanical entanglement, we consider three specific scenarios: (i) coherent coupling, (ii) dissipative coupling, and (iii) cooperative coupling (i.e., both the coherent and dissipative couplings are considered).
V.1 Coherent coupling ($g_{\kappa}=0$)
@@
In the coherent coupling case ($g_{\kappa}=0$), we present the entanglement measure $E_{N}$ as a function of both the coherent coupling strength $g_{\omega}$ and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$ in Fig. 3(a) to explore the conditions for optimal optomechanical entanglement at a bath temperature of $T=0.4$ K. The results reveal that the generated entanglement remains relatively weak compared to that typically observed in the conventional optomechanical system Vitali et al. (2007). This is because the stronger coupling strength will lead to system unstable. To numerically ensure system stable, the coupling strength is required to be weak, giving rise to weak entanglement. Nonetheless, optimal entanglement can still be achieved around the point ($\Delta_{s}/\omega_{m},g_{\omega}/2\pi)\approx(2,3.1)$, as marked by the green star, indicating that even under these constraints, entanglement can be optimized.
@@
Figure 3: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the coherent coupling strength $g_{\omega}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the coherent coupling strength $g_{\omega}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\omega}/2\pi=3.1$ Hz and $\Delta_{s}/\omega_{m}=2$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW.
@@
To further investigate the coherent coupling effect on entanglement, we plot $E_{N}$ as a function of $g_{\omega}$ in Fig. 3(b), where $\Delta_{s}\approx 2\omega_{m}$ is fixed. It is obviously shown that $E_{N}$ increases with $g_{\omega}$. Once up to its maximum, $E_{N}$ will have a sharp reduction by further increasing $g_{\omega}$. Such abrupt behavior corresponds to the transition from the stability to instability, as shown by the pink-shaded region. Additionally, when the bath temperature increases to $T=4$ K [see the blue dashed curve], the coupling for generation of quantum entanglement is required to be larger and $E_{N}$ has a significant reduction. This indicates that quantum entanglement can be affected by the bath temperature. To clearly evaluate the effect of bath temperature, we plot $E_{N}$ as a function of the bath temperature in Fig. 3(c), where $g_{\omega}/2\pi=3.1$ Hz is fixed to obtain optimal entanglement. As the temperature increases, $E_{N}$ gradually declines. When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, that is, quantum entanglment vanishes and the system becomes separable.
V.2 Dissipative coupling ($g_{\omega}=0$)
@@
In dissipative coupling case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the coherent coupling case, even surpassing the entanglement achieved in conventional optomechanical system (i.e., $E_{N}=0.35$Vitali et al. (2007). This is demonstrated in Fig. 4(a), where $E_{N}$ is plotted as a function of both the dissipative coupling strength ($g_{\kappa}$) and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$. The optimal entanglement is found at approximately ($\Delta_{s}/\omega_{m},g_{\kappa}/2\pi)\approx(0.1,20)$, as indicated by the green star. In contrast to the coherent coupling case [see Fig. 2(a)], where the optimal entanglement occurs near the mechanical red sideband at $\Delta_{s}/\omega_{m}=2$, the optimal entanglement in the dissipative coupling case is obtained when the cavity is nearly resonant with the driving field, specifically at $\Delta_{s}/\omega_{m}=0.1$, which is far removed from the mechanical red sideband.
Figure 4: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the dissipative coupling strength $g_{\kappa}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the dissipative coupling strength $g_{\kappa}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=0.1$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\kappa}/2\pi=19$ Hz and $\Delta_{s}/\omega_{m}=0.1$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW.
@@
Furthermore, in Fig. 4(b), we investigate the behavior of the optomechanical entanglement as a function of the dissipative coupling strength, while fixing $\Delta_{s}=0.1\omega_{m}$. The resulting behavior of $E_{N}$ mirrors that observed in the coherent coupling case (see the red curve). The abrupt drop in entanglement marks the transition of the system from a stable to an unstable region, as indicated by the pink area. Notably, a relatively large dissipative coupling strength is required to achieve the optimal entanglement. This suggests that in the regime dominated by dissipative coupling, stronger interactions are necessary to generate significant quantum effects. As the bath temperature increases, for instance to $T=4$ K, the entanglement is reduced (see the blue curve). To explore the influence of bath temperature on optomechanical entanglement, we fix $g_{\kappa}/2\pi=19$ Hz and present the results in Fig. 4(c). We observe a sharp decrease in entanglement as the system temperature increases from 0 K to 5 K. However, beyond this point, the entanglement becomes more tolerant to further temperature increases, suggesting that the dissipative coupling becomes more robust to thermal noise at higher temperatures. Remarkably, entanglement persists even up to 25 K, This indicates that strong, noise-resistant entanglement can be generated through dissipative optomechanical coupling, opening new possibilities for quantum information processing and communication in noisy
V.3 Cooperative coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)